OpenStudy (anonymous):

The area of the base of a prism is 50 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism?

OpenStudy (anonymous):

@iambatman @AnswerMyQuestions

OpenStudy (anonymous):

Sorry, no. There aren't any images, just the problem.

OpenStudy (anonymous):

Rectangular based prism Base shape: Rectangle, length 'L' and width 'W' Area of base: L × W Perimeter of base: 2(L+W) Surface area = 2LW + 2(L+W)H Circular based prism Base shape: Circle, radius 'R' Area of base: pR² Perimeter of base: 2pR Surface area = 2pR² + 2pRH Triangular based prism Base shape: Triangle: base 'b', height 'h', and sides S1, S2 and S3 Area of base: ½b×h Perimeter of base: S1+ S2 + S3 Surface area = bh + (S1+ S2 + S3)H

OpenStudy (anonymous):

Which one is it? :)

OpenStudy (anonymous):

I'm not sure.. It doesn't say what TYPE of prism it is.

OpenStudy (anonymous):

I'd go with Rectangular prism.

OpenStudy (anonymous):

Yeah, so would I. Can you answer it? (Explanation not required, as long as it isn't too long.)